OpenStudy (anonymous):

i dont get this! How many license plates can be made which have the form ##LLL? Anwsers: 1,757,600 175,600 3,840,500 2,947,000

OpenStudy (campbell_st):

so its a question about combinations.... are ## numbers or letters..? if letters its or a combination.... ## are numbers (0 to 9) and L are letters (A to Z) looks like 10 x 10 x 26 x 26 x26

OpenStudy (anonymous):

=#*#*L*L*L =10-from 0-9(number of digits) =26-from A-Z( number of letters) =10*10*26*26*26 =1,757,600

OpenStudy (anonymous):

ohh ok.. ya i got it :)